OpenStudy (dehelloo):

If someone gives you 42 x 29 can you easily do that and get the exact answer in 5 seconds or is it normal to just round and 40 x 30 and do the exact answer on paper?

OpenStudy (dehelloo):

As in can you 42 x 29 quickly and do all the rounding mentally?

OpenStudy (anonymous):

type 42 29 on the screen and then hit the Enter key. 1218 in 10 micro seconds on a mid 2010 IMac.

OpenStudy (anonymous):

42 * 10 ~~ 420 * 3 = 1260 - 42 = 1218. That's how I would look at it and that's how I do the math. It's pretty quick (~5 seconds).